OpenStudy (anonymous):

find a cubic equation whose roots are -4 and -6i

jimthompson5910 (jim_thompson5910):

if -6i is a root, then so is 6i since complex roots come in complex conjugate pairs

jimthompson5910 (jim_thompson5910):

so you really have these 3 roots: -4, -6i, 6i

OpenStudy (anonymous):

ok but how do you make that into an equation...its cubic because there are three roots correct?

jimthompson5910 (jim_thompson5910):

which means x = -4, x = -6i, x = 6i x+4 = 0, x^2 = 36i^2, or x^2 = 36i^2 x+4=0 or x^2 = -36 or x^2 = -36 x+4=0 or x^2 = -36 x+4=0 or x^2+36 = 0 (x+4)(x^2+36) = 0 keep going...

jimthompson5910 (jim_thompson5910):

and yes, all cubic equations have 3 complex roots

jimthompson5910 (jim_thompson5910):

any nth degree polynomial has n complex roots (fundamental theorem of algebra)

OpenStudy (anonymous):

ok I think I understand now. thank you

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

so is this the final answer? (x+4) (x+6)(x+6)=0

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

(x+4) (x+6)(x+6)=0 implies that x+6=0 ---> x =-6 is a root, but -6 is NOT a root

OpenStudy (tkhunny):

Too bad the question does not ask for what was provided. If we want the Complex Conjugate Pairs, the question MUST state "with real coefficients". Otherwise, (x+4)(x+6i)(x-2) = 0 would also be valid, or (x+4)(x+6i)^2 = 0 or infinitely many other options. Please do not try to factor x^2 + 36. That makes no sense.

OpenStudy (anonymous):

ok so just leave it as (x+4)(x^2+36)

jimthompson5910 (jim_thompson5910):

no you have to expand it out

OpenStudy (tkhunny):

Make sure the question says "with Real Coefficients". If it doesn't, you might want to have a chat with the author of the question.

jimthompson5910 (jim_thompson5910):

(x+4)(x^2+36) x(x^2+36) + 4(x^2+36) what's next?

OpenStudy (anonymous):

x^3+36x+4x^2+144=0 x^3-4x^2+36x+144=0

jimthompson5910 (jim_thompson5910):

you somehow turned +4x^2 into -4x^2

jimthompson5910 (jim_thompson5910):

otherwise, it's perfect

OpenStudy (anonymous):

sorry typo

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

so y = x^3+4x^2+36x+144 is your cubic equation

OpenStudy (anonymous):

oh ok it just clicked in my head. thanks I really appreciate your help

jimthompson5910 (jim_thompson5910):

ok great, yw